Use the Laplace transform to solve the following initial-value problem. y +5y +4y = 20 sin 2t, y(0)=-1, y(0) = 2

Note: Use partial fractions when solving
Answer #1

Let us assume the Laplace transform of y(t), L(y(t))=Y(s) then, we have L(y'(t))= sY(s)-y(0) and L(y"(t))= s​​​​​​2​​​​​Y(s)-sy(0)-y'(0) . By putting the initial values y(0)=-1 and y'(0)=2 and then taking the laplace transform both sides of the given differential equation, we will get the expression of Y(s).

Then, using the Method of Partial Fraction, we will get expression of Y(s) in partial fraction form. By taking the inverse Laplace transform on both sides of Y(s), and using the rule L​​​​-1​( s/(s​​​​​​2​​​​​ + b​2))= Cos (bt) and L​-1​(1/(s+a))= e​​​​​​-at

The step by step explanatory solution is provided below.

following gritial value problem. Y + sytly 20 cin at 22. Lly HJ = Y() solution use the Laplace transform to Lolve the ylo)Now , & Method of partial Fraction, Y(s) = (-83 38²- Us + 28) (x²+4) s²+55+ 4 33-362 us+ 28 8² + 4 4) (4+1) 4 + Ast 2²ty styPutting we B=0 and CF 7-UD in 5A+B+C+ 4D = -3 and A+C+D=-1, get 5A to +(7-400 +40.2-3. 3) 5A + +=-3 5 5A =-3-7 = 10. A=-2 OH- t § yut) = -2 Cos (2+) - cos (bt) ; (: LY ( PC -yt - 2 cos(at) - e +2 e s2+62 28 -at e LY ( _yt +(sta The solution is giv